A hairdresser mixes a 30% peroxide solution with a 10% peroxide solution to get 4 cups of20% peroxide solution. How much of the 10% peroxide solution did she use?
Answer by stanbon(75887)

A hairdresser mixes a 30% peroxide solution with a 10% peroxide solution to get 4 cups of20% peroxide solution. How much of the 10% peroxide solution did she use?
----------
Equation:
active + active = active
0.30x + 0.10(4-x) = 0.20*4
------
30x + 10*4 - 10x = 20*4
20x = 10*4
x = 2 cups (amt. of 30% solution needed)
---
4-x = 2 cups (amt. of 10% solution needed)
=====================
Cheers,
Stan H.
===============